How many liters of 50% acid must be mixed with 20% acid to get 10 liters of a 30% mixture?
Answer by Alan3354(69443)

How many liters of 50% acid must be mixed with 20% acid to get 10 liters of a 30% mixture?
-------------------
Do it like this:
if a 14% acid solution is mixed with a 19% acid solution to produce 45 gallons of a 15% acid solution how much (in gallons) of each solution should be used?
--------------
f = 14%
n = 19%
----
f + n = 45 (total gallons)
14f + 19n = 15*45 (total acid)
-----
n = 45 - f
Sub for n
14f + 19*(45-f) = 675
14f + 855 - 19f = 675
-5f = -180
f = 36 gallons of 14%
n = 9 gallons of 19%